One positive number is one-fifth of another number. The difference between the two numbers is 184, find the numbers.
Answer by stanbon(75887)

One positive number is one-fifth of another number. The difference between the two numbers is 184, find the numbers.
-----
x = (1/5)y
y - x = 184
--------
Substitute for "x" and solve for "y":
y - (1/5)y = 184
-----
(4/5)y = 184
------
y = (5/4)*184
-----
y = 5*46
----
y = 230
x = (1/5)230 = 46
=======
Cheers,
Stan H.
-------